# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v2

@@ 5.3 Examples and Results
1e-3 AIR $112.0e$ 11 $115.8e$ 11 $121.1e$ 12 $123.4e$ 12
MP-IR (1,1,1) $-13.446e\text{-}$ 1 $-179.7e\text{-}$ 17 $-13.632e\text{-}$ 1 $-161.5e\text{-}$ 16
(2,1,1) $-13.446e\text{-}$ 1 $-123.6e\text{-}$ 12 $-13.632e\text{-}$ 1 $-125.2e\text{-}$ 12
(2,2,1) $-13.446e\text{-}$ 1 $-126.1e\text{-}$ 12 $-13.632e\text{-}$ 1 $-121.7e\text{-}$ 12
(2,2,2) $-13.446e\text{-}$ 1 $-86.8e\text{-}$ 8 $-13.632e\text{-}$ 1 $-87.4e\text{-}$ 8
(3,2,1) $-13.446e\text{-}$ 1 $-61.6e\text{-}$ 6 $-13.632e\text{-}$ 1 $-79.3e\text{-}$ 7
(3,2,2) $-13.446e\text{-}$ 1 $-61.5e\text{-}$ 6 $-13.632e\text{-}$ 1 $-79.5e\text{-}$ 7
(3,3,2) $-13.445e\text{-}$ 1 $-51.7e\text{-}$ 5 $-13.633e\text{-}$ 1 $-67.6e\text{-}$ 6
(3,3,1) $-13.445e\text{-}$ 1 $-51.9e\text{-}$ 5 $-13.633e\text{-}$ 1 $-69.7e\text{-}$ 6
(3,3,3) $-13.442e\text{-}$ 1 $-47.5e\text{-}$ 4 $-13.634e\text{-}$ 1 $-46.6e\text{-}$ 4
1e-4 AIR $196.2e$ 19 $201.8e$ 20 $203.7e$ 20 $211.1e$ 21
MP-IR (1,1,1) $-12.129e\text{-}$ 1 $-163.4e\text{-}$ 16 $-13.289e\text{-}$ 1 $-162.7e\text{-}$ 16
(2,1,1) $-12.129e\text{-}$ 1 $-115.0e\text{-}$ 11 $-13.289e\text{-}$ 1 $-113.1e\text{-}$ 11
(2,2,1) $-12.129e\text{-}$ 1 $-116.8e\text{-}$ 11 $-13.289e\text{-}$ 1 $-111.4e\text{-}$ 11
(2,2,2) $-12.129e\text{-}$ 1 $-71.4e\text{-}$ 7 $-13.289e\text{-}$ 1 $-71.5e\text{-}$ 7
(3,2,1) $-12.129e\text{-}$ 1 $-52.2e\text{-}$ 5 $-13.289e\text{-}$ 1 $-54.0e\text{-}$ 5
(3,2,2) $-12.129e\text{-}$ 1 $-52.3e\text{-}$ 5 $-13.289e\text{-}$ 1 $-54.0e\text{-}$ 5
(3,3,2) $-12.128e\text{-}$ 1 $-41.3e\text{-}$ 4 $-13.289e\text{-}$ 1 $-41.1e\text{-}$ 4
(3,3,1) $-12.129e\text{-}$ 1 $-41.5e\text{-}$ 4 $-13.289e\text{-}$ 1 $-41.9e\text{-}$ 4
(3,3,3) $-12.124e\text{-}$ 1 $-31.5e\text{-}$ 3 $-13.277e\text{-}$ 1 $-31.4e\text{-}$ 3
Figure 3: Spectra example: graph of RRE vs. iteration number for the MP-IR results for various precision combinations for $3\%$ noise and $\alpha^{2}=1e\text{-}3$. Pane (a) focuses on precision combination (3, 3, 3)333(3,\,3,\,3)( 3 , 3 , 3 ) capturing the large variation in the yellow region compared to the other precision combinations. Pane (b) focuses on the variability of precision combinations (3, 3, 1)331(3,\,3,\,1)( 3 , 3 , 1 ) and (3, 3, 2)332(3,\,3,\,2)( 3 , 3 , 2 ), denoted in the blue region which is smaller than the yellow region in pane (a).
(Image Deblurring - reconstruction) - For our final example we consider the 2D image reconstruction problem which comes from the software package IR Tools [11]. Similarly to the reconstruction of Spectra, we compare the behavior of AIR against MP-IR for various precision combinations. The Hubble image for this problem in Figure 4a contains $512\times 512$ pixels which corresponds to a blurring operator of size $256^{2}\times 256^{2}$ making both the application of chop and the storage of the full problem untenable. However, because the point spread function (Figure 4b) is rank-1, the blurring matrix $A$ can be decomposed into a Kronecker product, each part of which is of reasonable size to compute its SVD in fp64, fp32, and fp16:
@@
With this Kronecker structure, it is possible to compute MP-IR efficiently; see [6] for more implementation details regarding chop and the Kronecker product. To efficiently implement AIR one may form a normal matrix $C$ that admits a spectral factorization by imposing periodic boundary conditions when convolving with the point spread function; see [16] for an overview of 2D image convolution. Because of the astronomical nature of the Hubble image, the choice of periodic or zero boundary conditions is typically experimentally equivalent. Zero boundary conditions were used in the formation of $A$. Figure 4c shows the blurred image with $1\%$ noise.
Figure 4: Image deblurring example: (a) true Hubble image ($512\times 512$ pixels), (b) PSF ($31\times 31$ pixels), (c) blurred and $1\%$ noised image ($256\times 256$ pixels).
Similarly to the numerical results shown for the Spectra reconstruction example, Table 4 provides the sRRE and standard deviation values at various precision combinations for MP-IR as well as AIR as a function of the regularization parameter and the noise level of the problem. Here, we note that we considered different noise levels: $1\%$ and $5\%$ and regularization parameters: $-11e\text{-}$ 1, $-21e\text{-}$ 2, and $-31e\text{-}$ 3 compared to the Spectra example, though again for the same reasons as before, we only show results for the two smaller regularization parameters. The same general behaviors observed for the Spectra example were also observed here with respect to observing larger sRRE standard deviations when utilizing a lower precision preconditioner.
We do note the observed rise in the sRRE value for the precision combination (3, 3, 3)333(3,\,3,\,3)( 3 , 3 , 3 ) involving Pr$=3$ 3 when the smaller $\alpha^{2}=1e\text{-}3$ was used. We attribute this behavior, in part, to the large variability experienced when using fp16 in the preconditioner (or equivalently, the solution basis) which can be seen through the standard deviation of the sRRE values in Table 4. From observing the effect that the choice of regularization parameter $-21e\text{-}$ 2 versus $-31e\text{-}$ 3 can make, it appears that this effect can be reduced with more regularization, but with some loss of accuracy. For conciseness, we do not show reconstructed images using AIR and MP-IR since the reconstructions amongst the latter for the various precision combinations are indiscernible and the former is meaningless given it’s RRE behavior.
Table 4: Image deblurring example: stable relative reconstructive error (i.e., the average RRE for iterations $3$ $10$) and the standard deviation for MP-IR and AIR methods as a function of the regularization parameter for $1\%$ and $5\%$ noise. MP-IR errors are provided for various mixed precision combinations.
@@
1e-2 AIR $31.9e$ 3 $34.6e$ 3 $39.2e$ 3 $42.3e$ 4
MP-IR (1,1,1) $-12.620e\text{-}$ 1 $-174.5e\text{-}$ 17 $-12.628e\text{-}$ 1 $-174.5e\text{-}$ 17
(2,1,1) $-12.620e\text{-}$ 1 $-87.2e\text{-}$ 8 $-12.628e\text{-}$ 1 $-85.7e\text{-}$ 8
(2,2,1) $-12.620e\text{-}$ 1 $-86.7e\text{-}$ 8 $-12.628e\text{-}$ 1 $-85.2e\text{-}$ 8
(2,2,2) $-12.620e\text{-}$ 1 $-86.0e\text{-}$ 8 $-12.628e\text{-}$ 1 $-85.5e\text{-}$ 8
(3,2,1) $-12.621e\text{-}$ 1 $-42.7e\text{-}$ 4 $-12.629e\text{-}$ 1 $-42.7e\text{-}$ 4
(3,2,2) $-12.621e\text{-}$ 1 $-42.7e\text{-}$ 4 $-12.629e\text{-}$ 1 $-42.7e\text{-}$ 4
(3,3,2) $-12.629e\text{-}$ 1 $-32.7e\text{-}$ 3 $-12.636e\text{-}$ 1 $-32.6e\text{-}$ 3
(3,3,1) $-12.629e\text{-}$ 1 $-32.7e\text{-}$ 3 $-12.636e\text{-}$ 1 $-32.6e\text{-}$ 3
(3,3,3) $-12.630e\text{-}$ 1 $-33.2e\text{-}$ 3 $-12.639e\text{-}$ 1 $-33.3e\text{-}$ 3
1e-3 AIR $113.9e$ 11 $121.4e$ 12 $121.8e$ 12 $125.2e$ 12
MP-IR (1,1,1) $-12.468e\text{-}$ 1 $-173.3e\text{-}$ 17 $-12.561e\text{-}$ 1 $-162.0e\text{-}$ 16
(2,1,1) $-12.468e\text{-}$ 1 $-71.2e\text{-}$ 7 $-12.561e\text{-}$ 1 $-88.6e\text{-}$ 8
(2,2,1) $-12.468e\text{-}$ 1 $-82.9e\text{-}$ 8 $-12.561e\text{-}$ 1 $-86.1e\text{-}$ 8
(2,2,2) $-12.468e\text{-}$ 1 $-71.0e\text{-}$ 7 $-12.561e\text{-}$ 1 $-87.7e\text{-}$ 8
(3,2,1) $-12.547e\text{-}$ 1 $-22.5e\text{-}$ 2 $-12.638e\text{-}$ 1 $-22.4e\text{-}$ 2
(3,2,2) $-12.547e\text{-}$ 1 $-22.5e\text{-}$ 2 $-12.638e\text{-}$ 1 $-22.4e\text{-}$ 2
(3,3,2) $-12.925e\text{-}$ 1 $-11.4e\text{-}$ 1 $-13.031e\text{-}$ 1 $-11.5e\text{-}$ 1
(3,3,1) $-12.925e\text{-}$ 1 $-11.4e\text{-}$ 1 $-13.031e\text{-}$ 1 $-11.5e\text{-}$ 1
(3,3,3) $-13.001e\text{-}$ 1 $-11.7e\text{-}$ 1 $-13.098e\text{-}$ 1 $-11.7e\text{-}$ 1
@@ 6 Conclusion
In this work, we have investigated IR on the Tikhonov problem in mixed precision. We have shown that the iterates of IR on the Tikhonov problem computed in mixed precision behave in a filtering manner by deriving methodology to formulate the iterates as a recursive relationship involving the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous terms. Our numerical results suggest that simulated mixed precision IR applied to inverse problems can give comparable accuracy compared to results computed in double precision and that solution variability when utilizing a low precision preconditioner can be influenced by the level of regularization as well as with the choice of the working and residual precisions. We also found that the MP-IR method to be superior when compared to the described AIR method which may be appropriate and efficient to compute with in image deblurring applications.
@@ Acknowledgments
The authors wish to thank the anonymous referees for carefully reading the manuscript and providing many useful suggestions that improved the presentation of the work. Additionally, they would like to thank Alan Edelman and Evelyne Ringoot for providing their invaluable computing expertise. L. O. would like to thank Marco Donatelli for hosting a research visit to Como in 2025 which included fruitful discussions that also helped improve the manuscript. This work was partially supported by the National Science Foundation (NSF) under grants DMS-2038118 and DMS-2208294.

### v3

@@ 5.3 Examples and Results
1e-3 AIR $2.0e11$ $5.8e11$ $1.1e12$ $3.4e12$
MP-IR (1,1,1) $3.446e\text{-}1$ $9.7e\text{-}17$ $3.632e\text{-}1$ $1.5e\text{-}16$
(2,1,1) $3.446e\text{-}1$ $3.6e\text{-}12$ $3.632e\text{-}1$ $5.2e\text{-}12$
(2,2,1) $3.446e\text{-}1$ $6.1e\text{-}12$ $3.632e\text{-}1$ $1.7e\text{-}12$
(2,2,2) $3.446e\text{-}1$ $6.8e\text{-}8$ $3.632e\text{-}1$ $7.4e\text{-}8$
(3,2,1) $3.446e\text{-}1$ $1.6e\text{-}6$ $3.632e\text{-}1$ $9.3e\text{-}7$
(3,2,2) $3.446e\text{-}1$ $1.5e\text{-}6$ $3.632e\text{-}1$ $9.5e\text{-}7$
(3,3,2) $3.445e\text{-}1$ $1.7e\text{-}5$ $3.633e\text{-}1$ $7.6e\text{-}6$
(3,3,1) $3.445e\text{-}1$ $1.9e\text{-}5$ $3.633e\text{-}1$ $9.7e\text{-}6$
(3,3,3) $3.442e\text{-}1$ $7.5e\text{-}4$ $3.634e\text{-}1$ $6.6e\text{-}4$
1e-4 AIR $6.2e19$ $1.8e20$ $3.7e20$ $1.1e21$
MP-IR (1,1,1) $2.129e\text{-}1$ $3.4e\text{-}16$ $3.289e\text{-}1$ $2.7e\text{-}16$
(2,1,1) $2.129e\text{-}1$ $5.0e\text{-}11$ $3.289e\text{-}1$ $3.1e\text{-}11$
(2,2,1) $2.129e\text{-}1$ $6.8e\text{-}11$ $3.289e\text{-}1$ $1.4e\text{-}11$
(2,2,2) $2.129e\text{-}1$ $1.4e\text{-}7$ $3.289e\text{-}1$ $1.5e\text{-}7$
(3,2,1) $129e\text{-}1 2.2$$2e\text{-}5$ $3.289e\text{-}1$ $4.0e\text{-}5$
(3,2,2) $2.129e\text{-}1$ $2.3e\text{-}5$ $3.289e\text{-}1$ $4.0e\text{-}5$
(3,3,2) $2.128e\text{-}1$ $1.3e\text{-}4$ $3.289e\text{-}1$ $1.1e\text{-}4$
(3,3,1) $2.129e\text{-}1$ $1.5e\text{-}4$ $3.289e\text{-}1$ $1.9e\text{-}4$
(3,3,3) $2.124e\text{-}1$ $1.5e\text{-}3$ $3.277e\text{-}1$ $1.4e\text{-}3$
Figure 3: Spectra example: graph of RRE vs. iteration number for the MP-IR results for various precision combinations for $3\%$ noise and $\alpha^{2}=1e\text{-}3$. Pane (a) focuses on precision combination (3, 3, 3)(3,\,3,\,3) capturing the large variation in the yellow region compared to the other precision combinations. Pane (b) focuses on the variability of precision combinations (3, 3, 1)(3,\,3,\,1) and (3, 3, 2)(3,\,3,\,2), denoted in the blue region which is smaller than the yellow region in pane (a).
(Image Deblurring - reconstruction) - For our final example we consider the 2D image reconstruction problem which comes from the software package IR Tools [11]. Similarly to the reconstruction of Spectra, we compare the behavior of AIR against MP-IR for various precision combinations. The Hubble image for this problem in Figure 4a contains $256\times 256$ pixels which corresponds to a blurring operator of size $256^{2}\times 256^{2}$ - making both the application of chop and the storage of the full problem untenable. However, because the point spread function (Figure 4b) is rank-1, the blurring matrix $A$ can be decomposed into a Kronecker product, each part of which is of reasonable size to compute its SVD in fp64, fp32, and fp16:
@@
With this Kronecker structure, it is possible to implement MP-IR efficiently; see [6] for more implementation details regarding chop and the Kronecker product. To efficiently implement AIR one may form a normal matrix $C$ that admits a spectral factorization by imposing periodic boundary conditions when convolving with the point spread function; see [16] for an overview of 2D image convolution. Because of the astronomical nature of the Hubble image, the choice of periodic or zero boundary conditions is typically experimentally equivalent. Zero boundary conditions were used in the formation of $A$. Figure 4c shows the blurred image with $1\%$ noise.
Figure 4: Image deblurring example: (a) true Hubble image ($256\times 256$ pixels), (b) PSF ($31\times 31$ pixels), (c) blurred and $1\%$ noised image ($256\times 256$ pixels).
Similarly to the numerical results shown for the Spectra reconstruction example, Table 4 provides the sRRE and standard deviation values at various precision combinations for MP-IR as well as AIR as a function of the regularization parameter and the noise level of the problem. Here, we note that we considered different noise levels: $1\%$ and $5\%$ and regularization parameters: $1e\text{-}1$, $1e\text{-}2$, and $1e\text{-}3$ compared to the Spectra example, though again for the same reasons as before, we only show results for the two smaller regularization parameters. The same general behaviors observed for the Spectra example were also observed here with respect to observing larger sRRE standard deviations when utilizing a lower precision preconditioner.
[…]
Table 4: Image deblurring example: stable relative reconstructive error (i.e., the average RRE for iterations 33 - $10$) and the standard deviation for MP-IR and AIR methods as a function of the regularization parameter for $1\%$ and $5\%$ noise. MP-IR errors are provided for various mixed precision combinations.
@@
1e-2 AIR $1.9e3$ $4.6e3$ $9.2e3$ $2.3e4$
MP-IR (1,1,1) $2.620e\text{-}1$ $4.5e\text{-}17$ $2.628e\text{-}1$ $4.5e\text{-}17$
(2,1,1) $2.620e\text{-}1$ $7.2e\text{-}8$ $2.628e\text{-}1$ $5.7e\text{-}8$
(2,2,1) $2.620e\text{-}1$ $6.7e\text{-}8$ $2.628e\text{-}1$ $5.2e\text{-}8$
(2,2,2) $2.620e\text{-}1$ $6.0e\text{-}8$ $2.628e\text{-}1$ $5.5e\text{-}8$
(3,2,1) $2.621e\text{-}1$ $2.7e\text{-}4$ $2.629e\text{-}1$ $2.7e\text{-}4$
(3,2,2) $2.621e\text{-}1$ $2.7e\text{-}4$ $2.629e\text{-}1$ $2.7e\text{-}4$
(3,3,2) $2.629e\text{-}1$ $2.7e\text{-}3$ $2.636e\text{-}1$ $2.6e\text{-}3$
(3,3,1) $2.629e\text{-}1$ $2.7e\text{-}3$ $2.636e\text{-}1$ $2.6e\text{-}3$
(3,3,3) $2.630e\text{-}1$ $3.2e\text{-}3$ $2.639e\text{-}1$ $3.3e\text{-}3$
1e-3 AIR $3.9e11$ $1.4e12$ $1.8e12$ $5.2e12$
MP-IR (1,1,1) $2.468e\text{-}1$ $3.3e\text{-}17$ $2.561e\text{-}1$ $2.0e\text{-}16$
(2,1,1) $2.468e\text{-}1$ $1.2e\text{-}7$ $2.561e\text{-}1$ $8.6e\text{-}8$
(2,2,1) $2.468e\text{-}1$ $2.9e\text{-}8$ $2.561e\text{-}1$ $6.1e\text{-}8$
(2,2,2) $2.468e\text{-}1$ $1.0e\text{-}7$ $2.561e\text{-}1$ $7.7e\text{-}8$
(3,2,1) $2.547e\text{-}1$ $2.5e\text{-}2$ $2.638e\text{-}1$ $2.4e\text{-}2$
(3,2,2) $2.547e\text{-}1$ $2.5e\text{-}2$ $2.638e\text{-}1$ $2.4e\text{-}2$
(3,3,2) $2.925e\text{-}1$ $1.4e\text{-}1$ $3.031e\text{-}1$ $1.5e\text{-}1$
(3,3,1) $2.925e\text{-}1$ $1.4e\text{-}1$ $3.031e\text{-}1$ $1.5e\text{-}1$
(3,3,3) $3.001e\text{-}1$ $1.7e\text{-}1$ $3.098e\text{-}1$ $1.7e\text{-}1$
@@ 6 Conclusion
In this work, we have investigated IR on the Tikhonov problem in mixed precision. We have shown that the iterates of IR on the Tikhonov problem computed in mixed precision behave in a filtering manner by deriving methodology to formulate the iterates as a recursive relationship involving the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous terms. Our numerical results suggest that simulated mixed precision IR applied to inverse problems can, on average, give comparable accuracy to within a few decimal places compared to results computed only in double precision. Additionally, our results suggest that the variability of the computed solution when utilizing a low precision preconditioner can be influenced by the level of regularization as well as by the choice of the working and residual precisions. We also found that the MP-IR method to be superior when compared to the described AIR method which may be appropriate and efficient to compute with in image deblurring applications.
@@ Acknowledgments
The authors wish to thank the anonymous referees and the handling editor for carefully reading the manuscript and providing many useful suggestions that improved the presentation of the work. Additionally, they would like to thank Alan Edelman and Evelyne Ringoot for providing their invaluable computing expertise. L. O. would like to thank Marco Donatelli for hosting a research visit to Como in 2025 which included fruitful discussions that also helped improve the manuscript. This work was partially supported by the National Science Foundation (NSF) under grants DMS-2038118 and DMS-2208294.